OpenStudy (anonymous):

Let's say that we have a square with all sides equal to 11 inches. What is the area of this square in square inches?

OpenStudy (anonymous):

121 inches squared. you just do 11 times 11.

OpenStudy (zehanz):

The area of a square with sides a is a²...

OpenStudy (anonymous):

A =s^2, s is the side. A = 11^2 = 121 inches^2

OpenStudy (anonymous):

thanks guys for the help.:)

OpenStudy (anonymous):

anytime.